On January 2nd 2016 lapar Corporation purchase a machine for $50,000 lapar paid shipping expensive $500 as well as installation cost of $1,200 the company estimated that the machine would have a useful life of 10 years and a residual value of 3000 on January 1st 2017 lapar made additional costing 36,000 to the machine in order to comply with Pollution Control ordinances. If lapar records depreciation expense under the straight-line method how much is the depreciation expense for 2017
Answer #1

Straight line depreciation = (cost of asset - salvage value)/ life of asset

Depreciation 2016 = (51,700 - 3,000) / 10 = $4,870

Book value as on 1 Jan 2017 = $51,700 - 4,870 = $46,830

Addition = $3,600

Depreciation 2017 = ($50,430 - 3,000)/9 = $5,270